Shrieves Casting Company is considering adding a new line to its product mix, and the capital budgeting analysis is being conducted by Sidney Johnson, a recently graduated MBA. The production line would be set up in unused space in the main plant. The machinery’s invoice price would be approximately $200,000, another $10,000 in shipping charges would be required, and it would cost an additional $30,000 to install the equip- ment. The machinery has an economic life of 4 years, and Shrieves has obtained a special tax ruling that places the equipment in the MACRS 3-year class. The machinery is expected to have a salvage value of $25,000 after 4 years of use. The new line would generate incremental sales of 1,250 units per year for 4 years at an incremental cost of $100 per unit in the first year, excluding depreciation. Each unit can be sold for $200 in the first year. The sales price and cost are both expected to increase by 3% per year due to inflation. Further, to handle the new line, the firm’s net working capital would have to increase by an amount equal to 12% of sales revenues. The firm’s tax rate is 40%, and its overall weighted average cost of capital, which is the risk-adjusted cost of capital for an average project (r), is 10%. a. Define “incremental cash flow.” (1) Should you subtract interest expense or dividends when calculating project cash flow? (2) Suppose the firm spent $100,000 last year to rehabilitate the production line site. Should this be included in the analysis? Explain. (3) Now assume the plant space could be leased out to another firm at $25,000 per year. Should this be included in the analysis? If so, how? (4) Finally, assume that the new product line is expected to decrease sales of the firm’s other lines by $50,000 per year. Should this be considered in the analysis? If so, how?
Answer #1
0 1 2 3 4
Incremental sales $ 2,50,000 $ 2,57,500 $ 2,65,225
Incremental cost $ 1,25,000 $ 1,28,750 $ 1,32,613
Depreciation $      79,992 $ 1,06,680 $     35,544 17784
Incremental NOI $      45,008 $      22,070 $     97,069
Tax at 40% $      18,003 $        8,828 $     38,827
NOPAT $      27,005 $      13,242 $     58,241
Add: Depreciation $      79,992 $ 1,06,680 $     35,544
Incremental OCF $ 1,06,997 $ 1,19,922 $     93,785
Capital expenditure = 200000+10000+30000 = $    2,40,000
Change in NWC $        30,000 $            900 $            927 $    -31,827
Salvage value of the machine $     25,000
Tax on gain = (25000-17784)*40% = $        2,886
Incremental project cash flow $ -2,70,000 $ 1,06,097 $ 1,18,995 $ 1,47,726
PVIF at 10% [PVIF = 1/1.1^n] 1 0.90909 0.82645 0.75131
PV at 10% $ -2,70,000 $      96,452 $      98,343 $ 1,10,989 $      35,783
NPV $        35,783
As the NPV is positive, the project can be implemented.
Incremental cash flow is the difference between:
*Cash flows with the project, and
*Cash flows without the project.
Yes, the 100000 to be spent at EOY 3, less the
tax on it, should be considered as a cash outflow in
the 3rd year. It should be discounted along with the
the net cash flow of the third year.
The after tax lease rent of 25000*(1-40%) = $15000
would be an opportunity cost and it should be
considered as a cash outflow.
The loss of sales of other products should also be
considered to the extent of the aftertax contribution
margin lost. It should be considered as a cash outflow.